OpenStudy (serenavdw):

Simplify 9(x-1)-2(x-)

OpenStudy (anonymous):

whats after x-?

OpenStudy (serenavdw):

oh sorry its (x-1)

OpenStudy (anonymous):

7x-7

OpenStudy (anonymous):

\[9 \times x \times 2 -4\]

OpenStudy (anonymous):

9(x-1)=? -2(x-1)=? 9(x-1)=9x-9 -2(x-1)=2x+2 9x-9-2x+2 =7x-7

OpenStudy (serenavdw):

thanks i got that to